# Higher-order homogenised riblet boundary conditions

### v1

@@ Appendix A Symmetry of the boundary-condition matrix
[…]
@@
$\left[\begin{matrix}\psi_{z}(y,L)\\ \psi(y,L)\end{matrix}\right]=\left[\begin{matrix}a&b\\ c&d\\ \end{matrix}\right]\left[\begin{matrix}\psi_{zz}(y,L)\\ -\psi_{zzz}(y,L)\end{matrix}\right],$
at distance $L$, the $[[a,b][c,d]]$ matrix will be symmetric. If this matrix is relayed back to $z=0$ using (4.7), it can be verified that it will still remain symmetric.
@@
$-\int_{\mathcal{D}}\overline{\varPhi_{11}}\mathrm{d}Y\mathrm{d}Z=\int_{\mathcal{D}}\left(\overline{V_{23}}_{,Y}+\overline{W_{23}}_{,Z}\right)\mathrm{d}Y\mathrm{d}Z=\int_{0}^{1}\overline{W_{23}}(Y,L)\mathrm{d}Y;$
@@
[…]
@@
Comparing (A.2) with (A.3) yields $g_{2}=h_{2}^{(p_{x})}$, consistently with what observed in table 1.
Four more symmetries were detected in table 1,
$f_{2}=-h_{2}^{(v_{xz})},\quad b_{3}=-c_{3},\quad h_{3}^{(p_{xy})}=-g_{3}^{(u_{xyz})},\quad f_{3}^{(p_{xx})}=-g_{3}^{(v_{xxz})},$
and will not be formally proved here. They have been, nonetheless, subjected to an additional numerical test using a riblet with smoothed edge, where numerical accuracy is much higher and allows them to be verified to several significant digits.
@@ Appendix B Proof that some coefficients turn out to be zero for any geometry
[…]
@@
$h_{2}=-\frac{1}{2\pi}\int_{\mathcal{D}}\frac{\partial}{\partial Y}\left(\overline{\varPhi_{11}}^{2}\right)\mathrm{d}Y\mathrm{d}Z=0$
because $\overline{\varPhi_{11}}$ (and thus also $\overline{\varPhi_{11}}^{2}$) is periodic in $Y$ of period $1$, and the integral of the derivative of a periodic function is always zero.
@@
[…]
@@
[…]
for the same reasons as in §B.1.
@@
$\begin{split}&h_{3}^{(nl)}=-\frac{1}{2\pi}\int_{\mathcal{D}}\varPhi_{1}\left(\overline{\varPsi_{21}}_{,Z}\overline{\varPhi_{11}}_{,Y}-\overline{\varPsi_{21}}_{,Y}\overline{\varPhi_{11}}_{,Z}\right)\mathrm{d}Y\mathrm{d}Z={}\\ &-\frac{1}{\pi}\int_{\mathcal{D}}\left[\overline{\varPsi_{21}}_{,Z}(\overline{\varPhi_{11}}^{2})_{Y}-\overline{\varPsi_{21}}_{,Y}(\overline{\varPhi_{11}}^{2})_{Z}\right]\mathrm{d}Y\mathrm{d}Z={}\\ &-\frac{1}{\pi}\int_{\mathcal{D}}\left(\overline{\varPsi_{21}}_{,Z}\overline{\varPhi_{11}}^{2}\right)_{Y}-\left(\overline{\varPsi_{21}}_{,Y}\overline{\varPhi_{11}}^{2}\right)_{Z}\mathrm{d}Y\mathrm{d}Z=0,\end{split}$
@@
$\begin{split}&e_{3}^{(nl)}=-\frac{1}{2\pi}\int_{\mathcal{D}}\overline{\varPsi_{21}}\left(\overline{\varPsi_{21}}_{,Z}\Delta_{2}\overline{\varPsi_{21}}_{,Y}-\overline{\varPsi_{21}}_{,Y}\Delta_{2}\overline{\varPsi_{21}}_{,Z}\right)\mathrm{d}Y\mathrm{d}Z={}\\ &-\frac{1}{\pi}\int_{\mathcal{D}}\left[(\overline{\varPsi_{21}}^{2})_{Z}\Delta_{2}\overline{\varPsi_{21}}_{,Y}-(\overline{\varPsi_{21}}^{2})_{Y}\Delta_{2}\overline{\varPsi_{21}}_{,Z}\right]\mathrm{d}Y\mathrm{d}Z={}\\ &-\frac{1}{\pi}\int_{\mathcal{D}}\left[(\overline{\varPsi_{21}}^{2})_{Z}\Delta_{2}\overline{\varPsi_{21}}\right]_{Y}-\left[(\overline{\varPsi_{21}}^{2})_{Y}\Delta_{2}\overline{\varPsi_{21}}\right]_{Z}\mathrm{d}Y\mathrm{d}Z=0,\end{split}$

### v2

@@ Appendix A Symmetry of the boundary-condition matrix
$\delta J=\int_{\mathcal{D}}\Delta_{2}\delta\psi\Delta_{2}\psi\mathrm{d}S=\int_{\mathcal{\partial D}}\left(\frac{\partial\delta\psi}{\partial n}\Delta_{2}\psi-\delta\psi\frac{\partial\Delta_{2}\psi}{\partial n}\right)\mathrm{d}l+\int_{\mathcal{D}}\delta\psi\underbrace{\Delta_{2}\Delta_{2}\psi}_{{}=\,0}\mathrm{d}S.$
@@
$\left[\begin{matrix}\psi_{z}(y,L)\\ \psi(y,L)\end{matrix}\right]=\left[\begin{matrix}a_{L}&-b_{L}\\ c_{L}&-d_{L}\\ \end{matrix}\right]\left[\begin{matrix}\psi_{zz}(y,L)\\ -\psi_{zzz}(y,L)\end{matrix}\right],$
at distance $L$ (notice the sign change of $b_{L}$ and $d_{L}$, with respect to (4.4), to compensate for the presence of $\psi_{zzz}$), the $a_{L},-b_{L}][c_{L},-d_{L}]]$ matrix will be symmetric. If this matrix is then relayed back to $z=0$ using (4.8) to produce $a_{1},-b_{2},c_{2},-d_{3}$ of (4.4), it can be verified that it will still remain symmetric, in particular we shall have $b_{2}=c_{2}$.
@@
$\int_{\mathcal{D}}\overline{\varPhi_{11}}\mathrm{d}Y\mathrm{d}Z=\int_{\mathcal{D}}\left(\overline{V_{23}}_{,Y}+\overline{W_{23}}_{,Z}\right)\mathrm{d}Y\mathrm{d}Z=\int_{0}^{1}\overline{W_{23}}(Y,L)\mathrm{d}Y;$
@@
$\int_{\mathcal{D}}\overline{\varPhi_{11}}\mathrm{d}Y\mathrm{d}Z=\int_{\mathcal{D}}\overline{\varPhi_{11}}\Delta_{2}\overline{U_{22}}\mathrm{d}Y\mathrm{d}Z=\int_{\partial\mathcal{D}}\left(\overline{\varPhi_{11}}\frac{\partial\overline{U_{22}}}{\partial n}-\frac{\partial\overline{\varPhi_{11}}}{\partial n}\overline{U_{22}}\right)\mathrm{d}l+\int_{\mathcal{D}}\underbrace{\Delta_{2}\overline{\varPhi_{11}}}_{=\,0}\overline{U_{22}}\mathrm{d}Y\mathrm{d}Z,$
@@
Comparing (A.2) with (A.3) yields $g_{2}=h_{2}^{(p_{x})}$, consistently with what is observed in table 1.
Four more identities were detected in table 1,
$f_{2}=-h_{2}^{(v_{xz})},\quad b_{3}=c_{3},\quad h_{3}^{(p_{xy})}=-g_{3}^{(u_{xyz})},\quad f_{3}^{(p_{xx})}=-g_{3}^{(v_{xxz})},$
all of which are only relevant to asymmetric riblets (the coefficients involved being zero for symmetric ones) and will not be formally proved here. They have, nonetheless, been subjected to an additional numerical test using a sawtooth riblet with smoothed edge, where numerical accuracy is much higher and has allowed them to be verified to several significant digits.
@@ Appendix B Proof that some coefficients turn out to be zero for any geometry
$h_{2}=-\frac{1}{2\pi}\int_{\mathcal{D}}\overline{\varPhi_{21}}^{\dagger}\,2\,\overline{\varPhi_{11}}_{,Y}\mathrm{d}Y\mathrm{d}Z,$
@@
$h_{2}=-\frac{1}{2\pi}\int_{\mathcal{D}}\frac{\partial}{\partial Y}\left(\overline{\varPhi_{11}}^{2}\right)\mathrm{d}Y\mathrm{d}Z$
and the latter integral is zero, because $\overline{\varPhi_{11}}$ (and thus also $\overline{\varPhi_{11}}^{2}$) is periodic in $Y$ of period $1$, and the integral of the derivative of a periodic function is always zero.
@@
$a_{2}=-\frac{1}{2\pi}\int_{\mathcal{D}}\overline{\varPsi_{31}}^{\dagger}\,4\,\Delta_{2}\overline{\varPsi_{21}}_{,Y}\mathrm{d}Y\mathrm{d}Z.$
@@
$a_{2}=-\frac{1}{2\pi}\int_{\mathcal{D}}\overline{\varPsi_{21}}\,4\,\Delta_{2}\overline{\varPsi_{21}}_{,Y}\mathrm{d}Y\mathrm{d}Z=\frac{2}{\pi}\int_{\mathcal{D}}\boldsymbol{\nabla}\overline{\varPsi_{21}}\boldsymbol{\cdot}\boldsymbol{\nabla}\overline{\varPsi_{21}}_{,Y}\mathrm{d}Y\mathrm{d}Z=\frac{1}{\pi}\int_{\mathcal{D}}\frac{\partial}{\partial Y}\left(\boldsymbol{\nabla}\overline{\varPsi_{21}}\right)^{2}\mathrm{d}Y\mathrm{d}Z,$
and the latter integral is zero for the same reason as in §B.1.
@@
$\begin{split}&h_{3}^{(nl)}=-\frac{1}{2\pi}\int_{\mathcal{D}}\left(\overline{\varPsi_{21}}_{,Z}\overline{\varPhi_{11}}_{,Y}-\overline{\varPsi_{21}}_{,Y}\overline{\varPhi_{11}}_{,Z}\right)\overline{\varPhi_{11}}\mathrm{d}Y\mathrm{d}Z={}\\ &-\frac{1}{\pi}\int_{\mathcal{D}}\left[\overline{\varPsi_{21}}_{,Z}(\overline{\varPhi_{11}}^{2})_{Y}-\overline{\varPsi_{21}}_{,Y}(\overline{\varPhi_{11}}^{2})_{Z}\right]\mathrm{d}Y\mathrm{d}Z={}\\ &-\frac{1}{\pi}\int_{\mathcal{D}}\left(\overline{\varPsi_{21}}_{,Z}\overline{\varPhi_{11}}^{2}\right)_{Y}-\left(\overline{\varPsi_{21}}_{,Y}\overline{\varPhi_{11}}^{2}\right)_{Z}\mathrm{d}Y\mathrm{d}Z=0,\end{split}$
@@
$\begin{split}&e_{3}^{(nl)}=-\frac{1}{2\pi}\int_{\mathcal{D}}\left(\overline{\varPsi_{21}}_{,Z}\Delta_{2}\overline{\varPsi_{21}}_{,Y}-\overline{\varPsi_{21}}_{,Y}\Delta_{2}\overline{\varPsi_{21}}_{,Z}\right)\overline{\varPsi_{21}}\mathrm{d}Y\mathrm{d}Z={}\\ &-\frac{1}{\pi}\int_{\mathcal{D}}\left[(\overline{\varPsi_{21}}^{2})_{Z}\Delta_{2}\overline{\varPsi_{21}}_{,Y}-(\overline{\varPsi_{21}}^{2})_{Y}\Delta_{2}\overline{\varPsi_{21}}_{,Z}\right]\mathrm{d}Y\mathrm{d}Z={}\\ &-\frac{1}{\pi}\int_{\mathcal{D}}\left[(\overline{\varPsi_{21}}^{2})_{Z}\Delta_{2}\overline{\varPsi_{21}}\right]_{Y}-\left[(\overline{\varPsi_{21}}^{2})_{Y}\Delta_{2}\overline{\varPsi_{21}}\right]_{Z}\mathrm{d}Y\mathrm{d}Z=0,\end{split}$